(no subject)

John Shields (jshields@ksera)
Thu, 27 May 1999 15:44:59 -0500

Shift Summary for Day shift, 05/27/99

On Shift: Wah
J. Shields

Background Accompaniment: most of the rest of the collaboration,
at one point or another.

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 8.5 hrs

------------------------------------------------

Accounting of lost time:

Rick Tesarek shut down mainframe 14 to evaluate HA HV problems. He
replaced 2 1443 HV cards and discovered that one of the spare 1443
cards was bad.

RK also worked on station 7 crate 51 (VV'). He removed ET board
latches in slots 5 and 6, since they've never been used. In theory,
this will make the crate more robust. cratemp.txt modified to
to skip those two modules.

SET moved delay jumper on crate 59 DYC from 8 to 2 microsec

Jason H. did a hot bit scan over 13K laser trigs, found ZERO hot bits.

Laser Q delayed another 2 nS. 8 nS of cable added, delay changed
in KTRIG by -6nS. Laser TDC should now read ~ 0.4 nS.

JAG fixed error in onsplit "cacher" script. Fixed COS cacher state
and restarted onsplitd. Seemed to have no errors.


========================================================================
Runs ended this shift:

13150 - random pulsar
13152 - cosmic run
13153 - hot bit scan

Tapes brought to FCC this shift: 0